# A dataset and model for recognition of audiologically relevant environments for hearing aids: AHEAD-DS and YAMNet+

### v3

@@ 6.3 Gain Sensitivity Analysis
Figure 7: The validation mAP and accuracy at different gain levels. The best results were achieved with zero gain. Higher mAP and accuracy indicate better results.
6.4 Model Speed and Edge Deployment
The mean processing times are shown in Figure 8. The tests were performed on a Google Pixel 3, a phone with modest specifications, released in 2018. We recorded 5, 10, 20, and 30s of audio on the phone, and measured the time it takes to process the audio by the model. For each length of time, we repeat the procedure five times and compute the mean time. We take a timestamp when the audio is buffered into memory and another timestamp when the results can be read from memory. The difference between the timestamps in milliseconds is recorded as the processing time. The processing time includes the time to compute the log mel spectrograms.
@@ 6.4 Model Speed and Edge Deployment
It takes a mean time of 200.4ms to process 5s of sound. There was approximately 50ms of latency to load the model, and approximately a linear increase of 30ms per 1 second of audio. In Figure 8, A straight line was fit to the data using linear regression. The line formula is $y=29.62x+48.31$. The Y-axis intercept is at 48.31. The line is almost a perfect fit to the data. This is a good indication that processing time increases linearly with recorded audio time. The latency of approximately 50ms is a one-off cost; so long as the variable representing the model remains in scope.
@@
This processing time indicates that it is feasible to create an application which is able to perform scene recognition in real time, as it takes significantly less time to process the audio compared to the audio playback time. There are far newer and more performant edge devices compared to our test system, so we expect lower latency and processing time with newer devices.
Figure 8: The mean processing time for 5, 10, 20, and 30s of sound over five iterations on a Google Pixel 3. It takes a mean time of 200.4ms to process 5s of sound. There was approximately 50ms of latency to load the model, and a linear increase of 30ms per 1 second of audio. A straight line was fit to the data using linear regression. The line formula is $y=29.62x+48.31$. The Y-axis intercept is at 48.31. The line is almost a perfect fit to the data. This is a good indication that processing time increases linearly with recorded audio time.
7 Discussion
7.1 Summary of Findings
The experiments demonstrate that YAMNet+ achieves promising performance, with a mAP of 0.83 and an accuracy of 0.93. The study highlights two critical findings that are essential for YAMNet+ to perform well. One is the necessity of transfer learning from the pre-trained YAMNet model on AudioSet to achieve satisfactory performance. The other is the necessity of proper adjustment of sound levels to ensure optimal performance.
The real-time sound-based scene recognition capabilities of YAMNet+ on edge devices were successfully demonstrated. On a Google Pixel 3 smartphone (a phone with modest specifications, released in 2018), the model processes audio with approximately 50ms of latency to load the model, and an approximate linear increase of 30ms per 1 second of audio. This indicates the feasibility of deploying such models for real-time applications on resource constrained devices.
@@ 7.1 Summary of Findings
The workflow and toolchain created for training and deploying a scene recognition model to edge devices is useful for future research in this area.
7.2 Comparison with Previous Work
@@
Compared to the original YAMNet public release, YAMNet+ is an improvement because we implemented a complete workflow, so it can be tuned with transfer learning. YAMNet published without training code can only be used as a black box. YAMNet+ is also lightweight with 3.7 million parameters compared to the >80>80 million parameters of other sound recognition models [17] [10].
AHEAD-DS allows comparison of models using a standardised dataset with consistent labels, divided into training/validation/testing sets, and annotated with the ground truth. AHEAD-DS is particularly valuable given the difficulty of directly comparing existing studies, where datasets may be incomplete, not in a ready-to-use state, labels differ between datasets, and the code and datasets may not be publicly available.
7.3 Limitations
The testing performed in previous sections reveal several limitations of the YAMNet+ and AHEAD-DS. Both training and validation loss of 0.53 show the AHEAD-DS dataset is small in size, and leads to underfitting when training YAMNet+ from scratch using random weight initialisation. We worked around this problem by using transfer learning, where the model is initialised using weights trained on AudioSet.
Another implication of the small dataset is the observed small improvement after applying regularisation techniques such as label smoothing and augmentation. A model must generalise better to perform well on a larger dataset, therefore regularisation techniques should be more impactful if the AHEAD-DS is expanded.
@@ 7.3 Limitations
The use of transfer learning comes with its own tradeoffs. We are locked to classifying 960 millisecond windows, as this is a hyperparameter which is fixed by the use of weights from an existing model. It is possible to work around the fixed length by adding the results from multiple windows and applying a softmax mathematical operation.
@@
The tests of gain sensitivity analysis indicate adjustment of the levels is necessary to achieve optimal performance. Recordings using new hardware configurations must be standardised, this only needs to be done once per device, to ensure recordings of conversations using the new device have similar levels to CHiME 6 Dev.
@@
The final notable limitation is the dependency on TensorFlow. YAMNet+ is built using the TensorFlow ML library, as it is necessary to maintain compatibility with YAMNet weights, required for transfer learning. TensorFlow has GPU compatibility problems, disabling the use of a GPU for quick model training and complicating the codebase.
7.4 Next Steps
@@
Future work can target three areas: the dataset, the model, and implementation details. The dataset is still relatively small compared to the likes of AudioSet and one finding is that transfer learning is essential to achieve reasonable performance. Therefore, the next step is to collect more data for a new version of AHEAD-DS. More sound clips, more varied sound environments, and more varied recording equipment configurations ensures models trained using a future version of AHEAD-DS are more generalisable.
@@
We can improve the dataset by choosing scene labels in a systematic way. In this paper we only adopted labels from an existing study and looked at several studies which apply labels to sound clips. We found a lack of rigorous studies to explain why these labels are good choices. Examples such as speech in noise seem very generic and one talker on subway platform seem oddly specific. We have yet to determine the balance between labels being too specific or too generic, concerning the acoustic environments and listening contexts critical from an audiological perspective. This is important as good scene labels are crucial in hearing research, especially when conducting Ecological Momentary Assessment (EMA) [12]. EMAs are common in hearing research and involve surveying participants in real-time, and scene recognition models can automate this process, saving time and reducing the participants’ cognitive load. However, if these labels are ambiguous, inconsistent, or too broad/narrow, model accuracy will suffer, negating the benefits of automation and potentially increasing the burden on participants.
@@
We may be able to improve the results by making better use of the limited data in AHEAD-DS. Speech and environment sounds can be left unmixed but standardised in the training set. More varied training data can be generated if they are mixed on-the-fly. So long as the test data is not mixed on-the-fly; results still remain comparable between test runs.
@@
Things that would improve the model implementation would include rebuilding the model by first cleanly separating feature computation and model code. This will enable further experimentation with other features such as time domain based features. This may enable more effective detection of acoustic properties such as reverberations, which are more salient in the time domain. Rebuilding the model enables the ability to test different architectures such as EfficientNet or ResNet. Newer architectures such as EfficientNet might boost precision and recall.
@@
We can improve the model implementation by using a modern ML library such as JAX or Pytorch. This allows us to avoid the complex legacy code and backward compatibility problems of TensorFlow. It would also significantly speed up training as more modern ML libraries have better compatibility with GPU acceleration.
We want to test the implementation on more types of edge hardware. This will allow us to determine the minimum resources necessary to enable real time sound recognition. If possible we want to deploy ML models to very power and resource constrained devices such as hearing devices and still maintain the ability to react to sounds in real time.
8 Conclusion
@@
This study addresses key challenges in scene recognition for hearing devices by introducing a refined dataset and machine learning model. AHEAD-DS is a publicly available dataset specifically designed for audiologically relevant environments, offering a standardised dataset for comparing models with consistent labels and in a ready-to-use state. Complementing this, YAMNet+ is presented as a robust sound recognition model, optimised for deployment on resource-constrained edge devices like smartphones. YAMNet+ only requires 3.7 million parameters compared to >80>80 million parameters of other sound recognition models. It serves as an open-source baseline, including model definitions, training, and testing code. Benchmarking YAMNet+ on AHEAD-DS yielded a mAP of 0.83 and an accuracy of 0.93. The model’s efficiency was demonstrated on a Google Pixel 3 smartphone, processing each second of audio in approximately 30ms with 50ms latency. The established workflow and toolchain for training and deploying such models on edge devices offer a valuable resource for future advancements in this field.
9 Acknowledgements
We thank our colleagues at the Macquarie University Hearing Hub and Google Research for their invaluable help, and the Google Australia Catalyst Fund and Macquarie University for their support.
9.1 Funding
@@
The authors disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work was supported by the Google DFI Catalyst fund and Macquarie University.
9.2 Declaration of conflicting interests
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
9.3 ORCID iDs
9.4 Links
Our website and code

### v4

@@ III-E Gain Sensitivity Analysis
Figure 5: The validation mAP and accuracy at different gain levels.
III-F Signal To Noise Ratio Analysis
The goal of these tests were to determine the impact of the mixing and level boosts outlined in Algorithm 1. The results are shown in Table 6.
In the original AHEAD-DS, a total of 2656 of the 6980 training files, 381 of the 1001 validation files, and 756 of the 1987 test files were processed using Algorithm 1. The default configuration of hyperparameters were used during training, while validation speech and environment sounds were mixed at a fixed SNRs rather than cycled through the range 10,5,0,5,10-10,-5,0,5,10 dB, both with and without the level boost.
Best performance was achieved when the validation parameters were closest to those used in training. A 0 dB SNR with a level boost achieved the best performance with mAP and accuracy at 0.84 and 0.91. Further drops in performance were observed when the SNR was further from 0 dB. There was a larger drop in performance without the level boost, at 0 dB SNR the mAP and accuracy were 0.80 and 0.87.
(a) SNR vs validation mAP. Higher values are better.
(b) SNR vs validation accuracy. Higher values are better.
Figure 6: The validation mAP, and accuracy for a series of SNRs.
III-G Best Per Class Threshold
The best per class thresholds on the validation set are shown in Table VI. The best threshold was calculated by finding the threshold which produced the largest value when precision is multiplied by recall.
TABLE VI: Best per class threshold on the validation set.
Class Threshold Precision Recall
cocktail_party 0.41 1.00 1.00
interfering_speakers 0.77 0.95 0.99
in_traffic 0.56 0.97 0.97
in_vehicle 0.58 0.98 0.99
music 0.66 0.90 0.99
quiet_indoors 0.35 1.00 1.00
reverberant_environment 0.21 0.97 0.99
wind_turbulence 0.79 0.99 0.99
speech_in_traffic 0.71 0.95 0.90
speech_in_vehicle 0.60 0.91 0.93
speech_in_music 0.93 0.98 0.85
speech_in_quiet_indoors 0.59 0.95 0.91
speech_in_reverb_env 0.75 0.93 0.90
speech_in_wind_turb 0.80 0.90 0.87
III-H Context Length
To extend the context beyond a single 960 millisecond window, results from a sequence of consecutive windows were aggregated. The aggregated results, as shown in Table 7, were evaluated on the validation set. Sequence lengths of 5, 10, 15, and 20 windows were tested. For aggregation, three methods were used: a simple moving average (SMA), a weighted moving average (WMA) with exponential weights, and a summation followed by a softmax function across classes. The exponential weights for the WMA were set as a series of powers of $e$, specifically $e^{0},e^{1},…,e^{n}$, where $n$ represents the number of windows in the sequence. Extending the context window beyond 960 millisecond windows improves performance. SMA and sum and softmax achieve the same accuracy as both methods are order-preserving functions, so the highest scoring class remains the same.
(a) Number of windows vs validation mAP. Higher values are better.
(b) Number of windows vs validation accuracy. Higher values are better. Simple moving average and sum and softmax produce the same accuracy.
Figure 7: Context length: The validation mAP and accuracy using a different number of windows.
III-I 16-Bit Model Performance and Speed on Edge Hardware
For edge computing deployment a 16-bit model was used. The 16-bit model achieved a validation mAP of 0.88 and an accuracy of 0.95. The processing times are shown in Table VII. The speed tests were performed on a Google Pixel 3, a phone with modest specifications, released in 2018. Segments of 5, 10, 20, and 30s of audio were recorded on the phone, and the time it takes to process the audio by the model was measured. For each length of time, this was repeated ten times.
Testing was not performed on streamed audio but on recorded audio. The demonstration application works by having a user press a button to start and then another to stop a recording. The application then buffers the last 5, 10, or 30 seconds of audio. To measure processing time, a timestamp was recorded when the audio is buffered into memory and when the results are available. The difference between these two timestamps, in milliseconds, is the total processing time, which includes the time required to compute the log mel spectrograms.
@@ III-H Context Length
It takes a mean time of 199.4ms to process 5s of sound with a SD of 10.31ms. There was approximately 50ms of latency to load the model, and approximately a linear increase of 30ms per 1 second of audio. A real-time factor (RTF) of between 0.038 and 0.044 was achieved over the ten iterations when processing 5s of sound. The RTF decreases with increasing length of processed audio. The latency of approximately 50ms is a one-off cost, so long as the variable representing the model remains in scope.
@@
TABLE VII: The processing times for 5, 10, 20, and 30s of sound over ten iterations on a Google Pixel 3.
Audio length (ms) Mean Processing Time (ms) $\pm$ SD RTF Range
5000 199.4 $\pm$ 10.31 0.037 to 0.044
10000 344.5 $\pm$ 12.29 0.032 to 0.036
20000 671.3 $\pm$ 22.88 0.032 to 0.035
30000 959.8 $\pm$ 31.12 0.030 to 0.033
III-J OpenYAMNet vs Perch vs AST
OpenYAMNet was compared against Perch and AST. The results are shown in Table VIII. The results were produced using the test set of AHEAD-DS. The two models were trained and validated on the embeddings from Perch and AST using the training and validation sets of AHEAD-DS respectively. Since both Perch and AST require an input length of 160000 (10 seconds), the results from OpenYAMNet were aggregated using the aforementioned sum and softmax procedure, across all 20 windows in each recording. Furthermore, because the models were only trained on the embeddings of Perch and AST, the version of OpenYAMNet used in this test had only the last layer trainable.
TABLE VIII: Comparison of OpenYAMNet against Perch and AST using AHEAD-DS.
Model Parameters Test mAP Test Accuracy
OpenYAMNet 3.7 million 0.84 0.93
Perch 7.8 million 0.80 0.89
AST 86.0 million 0.93 0.97
IV Discussion
This section discusses the results of the tests from the previous section. This includes the implications of the findings, a comparison to existing work, limitations of the current solution, and next steps for the future direction of the research.
IV-A Summary of Findings
@@
The experiments demonstrate that OpenYAMNet achieves promising performance, with a test mAP of 0.86 and an accuracy of 0.93. The workflow and toolchain created for training and deploying OpenYAMNet to edge devices is useful for future research in this area.
There were a few classes which were misidentified a notable number of times during testing. The class speech in music and music were often misidentified during pauses in speech and only music could be heard. The same issue occurs for speech in traffic and in traffic. The class speech in vehicle was sometimes misidentified as speech in wind turbulence because the sound of rushing air entering the vehicle can often be heard.
The study highlights two critical findings that are essential for OpenYAMNet to perform well. One is the necessity of transfer learning from the pre-trained YAMNet model on AudioSet to achieve good performance. The other is the necessity of proper adjustment of sound levels to ensure optimal performance.
Two ways of applying transfer learning were tested with OpenYAMNet. One method is to freeze all except the last layer, alternatively parameters in all layers are trainable. The Ablation Study showed that training all layers yielded better results than training only the last layer, which significantly impaired model performance. Another finding is that the TAU dataset is not a sufficient pre-training substitute for AudioSet. At approximately twelve thousand audio clips compared to the approximately two million in AudioSet, the TAU dataset is substantially smaller. The tradeoff of using transfer learning is that several hyperparameters become fixed. Changing a hyperparameter impacts all parameters dependent on the said hyperparameter. A new model will need to be initialised and re-trained from scratch using AudioSet, before transfer learning can be applied.
A learning rate of 0.001 was found to perform very poorly. The resulting ML model was likely not practically useful. This was likely because its large value caused the model parameters to overshoot optimal values.
The Ablation Study section tested several hyperparameter configurations. The two regularising techniques of label smoothing and augmentation did not have a major impact on the validation mAP and accuracy. The two techniques were kept enabled as the performance drop was minimal. The one noteworthy result was the validation loss of 0.01 when label smoothing was disabled. The result is expected as smoothing lowers the value of the true class while increasing the values of the negative classes during training. When a more extreme augmentation range using $\pm$24 dB was validated, there was a slight dip in performance. Since the gain is uniformly distributed between $\pm$24 dB, the distribution introduces occasional outliers at or near $\pm$24 dB leading to a performance drop. The use of weighted classes also caused a slight drop in performance. Focal loss already compensates for class imbalance. Therefore, the use of weighted classes does not seem necessary. AGC and normalisation caused a drop in performance. Both techniques discard information on the relative differences in level between environments.
A 2-fold cross validation test was conducted to determine if the previous experiment results were an artifact of the 70/10/20 training/validation/testing split. The 2-fold cross validation achieved equivalent mAP and accuracy to the test set. This is a strong indicator that the results produced using the 70/10/20 split are not an artifact of the split. The 70/10/20 split was based on a rule of thumb.
In gain sensitivity tests, deviations in sound levels outside the -5, 0, and 5 dB range led to diminished performance. This indicates adjustment of the levels is necessary to achieve optimal performance when applying the model to new datasets.
In SNR analysis there was a drop in performance without the level boost from Algorithm 1. This was due to a large drop in metrics for speech in quiet indoors. Without the level boost, the addition of speech did not significantly change the signal after quantisation. The reason to detect speech at -10 dB SNR was to test the possibility of detecting speech, though not intelligibility, in edge cases that would challenge a human with good hearing.
The validation results indicate it is possible to increase precision and accuracy by setting per class thresholds.
Extending beyond a single window smooths out the results, reducing the effects of occasional errors leading to overall higher mAP and accuracy. When using WMA with exponential weights and a context length of 20 windows, the mAP is lower than a single window and the accuracy also drops compared to 15 windows. Each 10 second audio sample is divided into 20 overlapping windows, the weight of the $20th$ window is $e^{20}$, which dominates the results of the previous 19 windows. If the $20th$ windows is misclassified, the entire audio sample will likely be misclassified, this is detrimental to performance. Sum and softmax show the largest increase to mAP. This is because softmax further boosts the score of the class with highest confidence relative to the scores of other classes.
The real-time sound-based scene recognition capabilities of OpenYAMNet on edge devices were successfully demonstrated on an edge device. With a RTF well below 1.0, this indicates the feasibility of deploying such models for real-time applications on resource constrained devices. There are more performant edge devices compared to our test system of a Google Pixel 3, even lower latency and processing time can be achieved with newer devices. The 16-Bit model deployed to the edge system achieved a 0.88 mAP, and 0.93 accuracy. The results indicate the 16-Bit model is practically indistinguishable from the 32-bit model.
IV-B Comparison with Previous Work
AHEAD-DS is based on HEAR-DS, which provides a solid foundation as the recordings of environment sounds have been open sourced with a permissive licence. One issue with HEAR-DS is that the dataset is not in a fixed, ready-to-use state. This is because some of the published data is unmixed, and some pseudo-random procedures were used during the production of the HEAR-DS dataset. Since the code has not been published, these procedures cannot be fully reproduced. The advantage of AHEAD-DS is that the dataset is in a ready-to-use state. This is important because it removes a variable when comparing ML models. The speech and environment sounds have been released in both mixed and unmixed forms, eliminating ambiguity.
OpenYAMNet was compared against Perch and AST. It outperforms Perch in both performance and efficiency by achieving a higher mAP, accuracy, while using fewer parameters, 0.84, 0.93, and 3.7 million respectively compared to Perch’s 0.80, 0.89, and 7.8 million. Although AST surpasses OpenYAMNet with the highest mAP and accuracy at 0.93 and 0.97, it is less efficient as it uses 86.0 million parameters.
@@ IV-B Comparison with Previous Work
Compared to the original YAMNet public release, OpenYAMNet is a fully open source re-implementation of YAMNet with a complete workflow, so it can be tuned with transfer learning. YAMNet was published without training code and can only be used as a black box.
IV-C Limitations
According to the authors of HEAR-DS, the 14 classes have been chosen because these classes are relevant to auditory scene recognition for hearing devices. These classes were chosen by audiologists and Audifon, a manufacturer of hearing aids. No further rationale was provided by the authors of HEAR-DS for these classes. Since AHEAD-DS adopts the same classes, this is a limitation inherited by AHEAD-DS.
Compared to datasets like AudioSet, AHEAD-DS is still relatively small, which is a disadvantage. As shown in the Ablation Study, training from scratch using AHEAD-DS leads to suboptimal results, scoring a validation loss of 0.53. Transfer learning is essential, as the model performs much better with a mAP of 0.86, when fine tuning weights pre-trained on AudioSet.
@@ IV-C Limitations
The use of transfer learning comes with its own tradeoffs. Hyperparameters are locked in due to transfer learning. This makes it resource intensive to change hyperparameters such as the window length. Extending it beyond 960ms will require the model be trained from scratch using AudioSet. It is possible to work around the fixed length by aggregating the results from multiple windows and applying a softmax operation.
@@
The tests of gain sensitivity and SNR analysis indicate adjustment of the levels is necessary to achieve optimal performance. Recordings using new hardware configurations must be standardised, this only needs to be done once per device, to ensure recordings of conversations using the new device have similar levels to CHiME 6 Dev.
@@
A notable limitation is the dependency on TensorFlow. OpenYAMNet is built using the TensorFlow ML library, as it is necessary to maintain compatibility with YAMNet weights, required for transfer learning. TensorFlow has GPU compatibility problems, disabling the use of a GPU for quick model training and complicating the codebase.
IV-D Next Steps
@@
Future work can target three areas: the dataset, the model, and implementation details. The dataset is still relatively small compared to the likes of AudioSet. Therefore, the next step is to collect more data for a new version of AHEAD-DS. More sound clips, more varied sound environments, and more varied recording equipment configurations ensures models trained using a future version of AHEAD-DS are more generalisable.
@@
The dataset can be improved by choosing scene labels in a systematic way. In this paper class labels are adopted labels from an existing study. There is a lack of rigorous studies to explain why these labels are good choices. Examples such as speech in noise seem very generic and one talker on subway platform seem oddly specific. The balance between labels being too specific or too generic are yet to be determined. This is important as good scene labels are crucial in hearing research, especially when conducting Ecological Momentary Assessment (EMA) [13]. EMAs are common in hearing research and involve surveying participants in real-time, and scene recognition models can automate this process, saving time and reducing the participants’ cognitive load. However, if these labels are ambiguous, inconsistent, or too broad/narrow, model accuracy will suffer, negating the benefits of automation and potentially increasing the burden on participants.
@@
It may be possible to improve the results by making better use of the limited data in AHEAD-DS. Speech and environment sounds can be left unmixed but standardised in the training set. More varied training data can be generated if mixed on-the-fly. So long as the test data is not mixed on-the-fly, results still remain comparable between test runs.
It may be possible to achieve better results by using strong annotations with AHEAD-DS. Unlike weak annotations, strong annotations precisely mark the start and end times of each sound.
@@
Things that would improve the model implementation include rebuilding the model by first cleanly separating feature computation and model code. This will enable further experimentation with other features such as time domain based features. This may enable more effective detection of acoustic properties such as reverberations, which are more salient in the time domain. Rebuilding the model enables the ability to test different architectures such as EfficientNet or ResNet. Newer architectures such as EfficientNet might boost precision and recall.
@@
The model implementation can be improved by using a modern ML library such as JAX or PyTorch. This allows us to avoid the complex legacy code and backward compatibility problems of TensorFlow. It would also significantly speed up training as more modern ML libraries have better compatibility with GPU acceleration.
A battery and thermal impact test of the ML model on various types of edge hardware would be a valuable contribution. This will help determine the feasibility of using the model on more types of highly resource constrained devices such as hearing devices.
V Conclusion
@@
This study addresses key challenges in scene recognition for hearing devices by introducing a refined dataset and machine learning model. AHEAD-DS is a publicly available dataset repacked and refined from several other open source datasets. It is designed for auditory scene recognition for hearing devices and offers a standardised dataset for comparing models with consistent labels and in a ready-to-use state. Complementing this, OpenYAMNet is presented as a robust sound recognition model, it only requires 3.7 million parameters, and optimised for deployment on resource-constrained edge devices like smartphones. It serves as an open-source baseline, including model definitions, training, and testing code. Benchmarking OpenYAMNet on AHEAD-DS yielded a mAP of 0.86 and an accuracy of 0.93. The model’s efficiency was demonstrated on a Google Pixel 3 smartphone, processing each second of audio in approximately 30ms with 50ms latency. The established workflow and toolchain for training and deploying such models on edge devices offer a valuable resource for future advancements in this field.
Acknowledgments
@@
This work was supported by the Google DFI Catalyst fund and Macquarie University.